OpenStudy (anonymous):

What is the equation of the line that passes through (-2, 2) and (1, -4)? a. y = 2x- 2 b. y = -2x- 4 c. y = -2x- 2 d. y = 2x+ 4

OpenStudy (mertsj):

Find the slope first.

OpenStudy (yrelhan4):

An easier method... put the points in each option and see which one satisfies.

OpenStudy (texaschic101):

set 1 (-2,2)....set 2 (1,-4) m = (y2 - y1) / (x2 - x1) -- m is the slope m = (-4 -2) / (1 - (-2) m = -6/3 = - 2 slope is -2 y = mx + b using points (1,-4) -4 = -2(1) + b -4 = - 2 + b - 2 = b (b is the y intercept) ANSWER : y = -2x - 2